if random variable is uniformly distributed (-2,1) find pdf of y = 2x ^3

1
Expert's answer
2022-01-28T14:03:05-0500

Since the random variable "X" is uniformly distributed with the interval, (-2,1), its pdf is given as,

"f(x)={1\over b-a}={1\over 1--2}={1\over3}"  

Therefore, the probability density function for the random variable "X" IS,

"f(x)={1\over3},\space -2\leq x\leq 1\\0,\space elsewhere"

To determine the pdf of "y=2x^3", we shall apply the cumulative density function (CDF) method as described below.

 We determine "G(y)," the  "cdf" of "y=2x^3",

Now,

"G(y)=p(Y\leq y)=p(2x^3\leq y)\implies p(x\leq({y\over2})^{1\over3})"

From definition of probability for continuous distributions,

 "p(x\leq({y\over2})^{1\over3})=\displaystyle\int^{({y\over2})^{1\over3}}_{-2}f(x)dx"

So,

"G(y)=p(x\leq({y\over2})^{1\over3})=\displaystyle\int^{({y\over2})^{1\over3}}_{-2}{1\over3}dx={x\over3}|^ {({y\over2})^{1\over3}} _{-2}= {{({y\over2})^{1\over3}}\over3}+{2\over3}"

To find the "pdf" of the random variable "y", we differentiate "G(y)" with respect to "y". That is,

"g(y)={dG(y)\over dy}={{1\over3}\times{1\over2}\times({y\over2})^{-{2\over3}}\over3}={1\over18}({y\over2})^{-{2\over3}}"

The limits are,

Lower limit,

"y=2\times(-2)^3=-16"

Upper limit,

"y=2\times(1)^3=2"

Therefore, the probability density function of the random variable "y"  is,

 "g(y)={1\over18}({y\over2})^{-{2\over3}},\space -16\leq y\leq 2\\0,\space elsewhere"